If the grandfather clock is running slow, how would you adjust the length of the pendulum to correct the timer?


1
Expert's answer
2021-12-02T10:07:41-0500

Since the period of oscillations (time to make one oscillation) is proportional to a square root of length, you need to decrease the length:


"T=2\pi\sqrt{\frac lg}"